The compound interest in 2years is 105 RS and the compound interest in 3years is 169 RS then find the rate of interest

Grade:12th pass

1 Answers

Swarnendu Chakraborty
6 years ago
p(1+r/100)²= 105+pp(1+r/100)³=169+psolve it to get r.. trick is simple..p will get eliminated...got It ???